OpenStudy (anonymous):

Finals are over, and you are moving back home for the summer. you need to rent a truck to move your possessions from the college residence hall to your home. You contact two local rental companies and acquire the following information for the 1 day cost of renting a truck. Company 1: $24.90 per day, plus $0.41per mile Company 2: $18.95 Per day, plus $0.26 per mile let x represent the number of miles driven in one day and C the total daily rental cost ($).

OpenStudy (anonymous):

Whats the question?

OpenStudy (anonymous):

I have to figure out what the costs are associated with each company to move I have already figured out the symbolic rules for both which are company A: $0.41x+ 24.90 Company B: $0.67x +18.95 The equation that I need to figure out is based off of 0.41x+24.90=0.67+18.95 this is where I am stuck because I have to figure out the miles

OpenStudy (precal):

just solve the equations you listed above, but you forgot your x in the right side next to .67x once you find x then that is the number of miles where both companies charge they same price

OpenStudy (anonymous):

Still not having any luck used the number of miles the example gave me of 18 but no luck

OpenStudy (anonymous):

subtract .67 from.41

OpenStudy (anonymous):

-.26x+24.9=18.95' subtract 24.9from 18.95 -0.26x=-5.95 whats 5.95 divided by 0.26?

OpenStudy (anonymous):

22.88or rounded up 24.9

OpenStudy (anonymous):

meant 22.9

OpenStudy (anonymous):

ok I still have to figure out if company 1 or 2 has the lowest price

OpenStudy (precal):

this is a system of equations. Your mileage will determine which company has the lowest price.